# Dimension-independent convergence rates of randomized nets using median-of-means

### v1

@@ 5 Tractability under more general randomization
$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}2^{m-\tilde{t}_{u}}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{(1-\theta^{\prime})dm/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
@@
$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{(1-\theta^{\prime})dm/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
@@
$\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}\Pr(Z(\boldsymbol{k})=1)|\hat{f}_{u}(\boldsymbol{k})|^{2}\leqslant 2^{-m+\tilde{t}_{u}}\frac{\|f\|^{2}_{u,\alpha,\lambda}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}$
for $T_{u,m}$ defined by equation (39). From Equation (9) and (16),
$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}\leqslant\|f\|^{2}_{\alpha,\lambda}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{-m+\tilde{t}_{u}}\frac{\gamma^{2}_{u}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$ (40)
Our conclusion finally follows from Lemma 2 and equation (5). ∎
@@
Consider again settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$. We assume equation (34) holds for weights $\{\gamma_{j},j\in 1{:}s\}$. Setting $\theta^{\prime}=1-\eta/d$ for $\eta>0$, the analysis in Remark 3 implies $\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\leqslant C_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for $C_{\eta}$ independent of $s$. Therefore, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent error bounds if $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ remains bounded as $s\to\infty$, which depends on growth of $\tilde{t}_{u}$.
@@
for an absolute constant $C$. For simplicity, we continue our discussion with the assumption $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for a nonnegative sequence $\{\tilde{t}_{j},j\in 1{:}s\}$. We can then bound
@@
When $d<\alpha+\lambda$, we can refine our argument to establish strong tractability under a weaker condition analogous to that outlined in Remark 7.
@@
for a constant $\Lambda$ independent of $s$. Let $\delta^{\prime}=\min(\Lambda^{-\phi},1/16)$, $\theta^{\prime}=1-\eta d,$
$\Gamma(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}$
@@
Therefore, $T_{u,m}\leqslant(\alpha+\lambda)(m-\tilde{t}_{u})$ and an argument similar to Corollary 2 shows
@@
$\leqslant$ $\frac{\delta^{\prime}}{\Gamma(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)},$
@@
$\Gamma^{\prime}=\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{2-\frac{2\phi\theta(\alpha+\lambda-d)}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|}2^{\tilde{t}_{u}\big{(}1+2\theta(\alpha+\lambda)-2\theta\phi(\alpha+\lambda-d)\big{)}}.$
Like in Remark 6, we can show $\Gamma(m)\leqslant C^{\prime}_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C^{\prime}_{\eta}$ independent of $s$ when equation (42) holds. Furthermore, we notice that $\alpha+\lambda>\phi(\alpha+\lambda-d)$ and
@@
which is bounded as $s\to\infty$ given equation (42). Our conclusion then follows from Lemma 2 with $\delta=1/16$ and equation (5). ∎
@@
For random linear scrambling using base-2 Niederreiter sequences, substituting the bound (41) into equation (42) yields the simplified tractability condition:
@@
Since $\phi>1$, this condition is strictly weaker than the requirements in [8, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, strong tractability holds if
@@ 6 Numerical experiments
In this section, we test our theory across integrands with varying dimension and decay on $\gamma_{u}$. The test integrands have the common form
@@
where $s=10,100,1000$ controls the dimension of integration domain, $\gamma=2,3,4$ controls the decay of $\gamma_{u}$ and $c_{\alpha}$ for $\alpha=0,1$ is a normalizing constant to ensure $\|f\|_{\alpha,1}=1$. To determine $c_{\alpha}$, we start by computing the ANOVA component given by equation (8):
@@
Hence,
$V^{(|u|)}_{1}(f_{u})=\|f^{(1,\dots,1)}_{u}\|_{L^{2}}=\frac{c^{|u|}_{\alpha}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}.$
By equation (12), for $v\subseteq u$,
$\Big{(}V^{v}_{\lambda}(f_{u})\Big{)}^{2}$
$\leqslant$ $\frac{c^{2|u|}_{\alpha}}{(\prod_{j\in u}j)^{2\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|v|}\Big{(}\int_{0}^{1}(x\exp(x)-1)^{2}\,\mathrm{d}x\Big{)}^{|u|-|v|}$
$\leqslant$ $\Big{(}V^{|u|}_{\lambda}(f_{u})\Big{)}^{2},$
while $V^{u}_{\lambda}(f_{u})=V^{|u|}_{\lambda}(f_{u})$. Therefore,
$\|f\|_{u,0,1}=\sup_{v\subseteq u}V^{v}_{1}(f_{u})=V^{u}_{1}(f_{u})=\frac{c^{|u|}_{\alpha}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}$
@@
A similar calculation shows $\|f_{s,\gamma,1}\|_{1,1}=1$ if we set
@@
With the above choice of $c_{\alpha}$, it also follows that $\gamma_{u}=\prod_{j\in}\gamma_{j}$ for $\gamma_{j}=j^{-\gamma}$.
For each integrand $f_{s,\gamma,\alpha}$, we evaluate the performance of the median estimator $\hat{\mu}^{(r)}_{E}$ for $m=1,2,\dots,16$ under two randomization schemes: the completely random design (CRD) and the random linear scrambling (RLS). The estimator $\hat{\mu}^{(r)}_{E}$ is computed as the sample median of $12r-$ 1 independent realizations of $\hat{\mu}_{E}$, where $\hat{\mu}_{E}$ is defined as in equation (2) with precision $E=64$. The observed integration errors in our tested range of $m$ is much larger than $2^{-E}$, ensuring that $\hat{\mu}^{(r)}_{E}$ closely approximates $\hat{\mu}^{(r)}_{\infty}$. Consequently, the error bounds in Theorem 1, 2 and 3 apply to $\hat{\mu}^{(r)}_{E}$ with probability at least $1-2^{-r}$. We set $r=10$, limiting the failure probability per data point to below $0.1\%$. As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the average of the $12r-$ 1 $\hat{\mu}_{E}$ realizations used by RLS. The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [11].
@@
Figure 1: Integration error for $f_{s,\gamma,\alpha}$ with $\alpha=0$.
@@
Figure 2: Integration error for $f_{s,\gamma,\alpha}$ with $\alpha=1$.
@@
$\gamma^{\prime}_{u}=\frac{\|f_{s,\gamma,0}\|_{1,1,u}}{\|f_{s,\gamma,0}\|_{1,1}}\leqslant\prod_{j\in u}\Big{(}\frac{c_{0}}{c_{1}}j^{-\gamma}\Big{)}=\prod_{j\in u}\gamma^{\prime}_{j}$
for $\gamma^{\prime}_{j}=(c_{0}/c_{1})j^{-\gamma}$. In general, this analysis extends to arbitrary norms $\|\cdot\|_{\alpha^{\prime},\lambda^{\prime}}$, $\alpha^{\prime}\in\mathbb{N},\lambda^{\prime}\in(0,1]$, and the corresponding $\gamma_{j}$ is always proportional to $j^{-\gamma}$. Thus, the decay rate $\gamma$ dominates convergence behavior as $s\to\infty$, irrespective of $\alpha$.
Next, we observe that for a fixed value of $\gamma$, the integration errors exhibit similar behavior as $s$ varies, aligning with our analysis of dimension-independent convergence. Applying equation (43) with parameters $d=1$, $p=\gamma$ and $q=1+\eta$ for any $\eta>0$, our theory predicts that for large $s$, $\hat{\mu}^{(r)}_{\infty}$ under RLS achieve a near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rate, provided $\alpha+\lambda+1/2<(\gamma+1)/2$. Empirically, the observed rates approximate $O(2^{-(3/2)m})$ when $\gamma=2$ and $O(2^{-2m})$ when $\gamma=3$. For $\gamma=4$, the convergence rates slightly exceed $O(2^{-2m})$ but remain significantly slower than $O(2^{-(5/2)m})$. This discrepancy is likely due to the limited range of $m$ tested. Notably, while our theory suggests that $\hat{\mu}^{(r)}_{\infty}$ under CRD can attain a near $O(2^{-(\alpha+\lambda+1/2)m})$ rate for $\alpha+\lambda+1/2<\gamma$, practical results show that CRD consistently underperforms RLS. We hypothesize that the superior theoretical rates for CRD may only manifest at substantially larger $m$.
Finally, we compare the integration errors under RLS and STD. Across the vast majority of test cases, RLS demonstrates superior performance over STD, reinforcing that the median RQMC approach (via RLS) is a more robust alternative to traditional RQMC (via STD) in high-dimensional settings. We also note that STD exhibits convergence rates comparable to those of RLS and substantially surpasses the theoretical $O(2^{-(3/2)m}$ rate derived from the root mean squared error when $\gamma=3$ or $4$. This discrepancy arises because each estimator $\hat{\mu}_{\infty}$ in STD is a special case of $\hat{\mu}^{(r)}_{\infty}$ with $r=1$, which satisfies the rapidly converging probabilistic error bounds in Theorems 2 and 3. However, averaging multiple replicates in STD compromises these bounds, as even a single outlier among the replicates can dominate the error and inflate its magnitude. This is the key mechanism that motivates our shift toward median RQMC approaches.

### v2

@@ 5 Tractability under more general randomization
$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}2^{m-\tilde{t}_{u}}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
@@
$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
@@
$\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}\Pr(Z(\boldsymbol{k})=1)|\hat{f}_{u}(\boldsymbol{k})|^{2}\leqslant 2^{-m+\tilde{t}_{u}}\frac{\|f\|^{2}_{u,\alpha,\lambda}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$
From Equation (9) and (16),
$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}=$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}\Pr(Z(\boldsymbol{k})=1)|\hat{f}_{u}(\boldsymbol{k})|^{2}$ (41)
$\leqslant$ $\|f\|^{2}_{\alpha,\lambda}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{-m+\tilde{t}_{u}}\frac{\gamma^{2}_{u}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$
Our conclusion follows from Lemma 2 and equation (5) after plugging in $T_{u,m}$ from equation (40). ∎
@@
Consider again settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$. We assume equation (34) holds for weights $\{\gamma_{j},j\in 1{:}s\}$. Setting $\theta^{\prime}=1-\eta/d$ for $\eta>0$, the analysis in Remark 3 implies $\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\leqslant C_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C_{\eta}$ independent of $s$. Therefore, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent error bounds if $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ remains bounded as $s\to\infty$, which depends on growth of $\tilde{t}_{u}$.
@@
for an absolute constant $C$.
For simplicity, we continue our discussion with the assumption $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for nonnegative numbers $\{\tilde{t}_{j},j\in 1{:}s\}$. We can then bound
@@
When $d<\alpha+\lambda$, we can refine our argument to establish dimension-independent convergence under a weaker condition analogous to that outlined in Remark 7.
@@
for a constant $\Lambda$ independent of $s$. Let $\delta^{\prime}=\min(\Lambda^{-\phi},1/16)$, $\theta^{\prime}=1-\eta/d,$
$\Gamma(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}$
@@
Therefore, $T_{u,m}\leqslant(\alpha+\lambda)(m-\tilde{t}_{u})$ and an argument similar to that used in Corollary 2 shows
@@
$\leqslant$ $\frac{\delta^{\prime}}{\Gamma(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)},$
@@
$\Gamma^{\prime}=\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{2-\frac{2\theta\phi(\alpha+\lambda-d)}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|}2^{\tilde{t}_{u}\big{(}1+2\theta(\alpha+\lambda)-2\theta\phi(\alpha+\lambda-d)\big{)}}.$
Analogous to Remark 6, we can show $\Gamma(m)\leqslant C^{\prime}_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C^{\prime}_{\eta}$ independent of $s$ when equation (43) holds. Furthermore, we notice that $\alpha+\lambda>\phi(\alpha+\lambda-d)$ and
@@
which is bounded as $s\to\infty$ because of equation (43). Our conclusion then follows from Lemma 2 with $\delta=1/16$ and equation (5). ∎
@@
For the random linear scrambling using base-2 Niederreiter sequences, substituting the bound (42) into equation (43) yields the simplified tractability condition:
@@
Since $\phi>1$, this condition is strictly weaker than the requirements in [9, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, $\mu^{(r)}_{\infty}$ attains dimension-independent convergence if
@@ 6 Numerical experiments
In this section, we test our theory across integrands with varying dimensions and decay rates on $\gamma_{u}$. The test integrands have the common form
@@
where $s=10,100,1000$ controls the dimension of integration domain, $\gamma=2,3,4$ controls the decay rate of $\gamma_{u}$ and $c_{\alpha}$ for $\alpha=0,1$ is a normalizing constant to ensure $\|f\|_{\alpha,1}=1$. To determine $c_{\alpha}$, we start by computing the ANOVA component given by equation (8):
@@
Hence, $V^{(|u|)}_{1}(f_{u})=\|f^{(1,\dots,1)}_{u}\|_{L^{2}}$. As discussed in Section 3, $\|f\|_{u,\alpha,\lambda}=V^{u}_{\lambda}(f_{u})$ when $\alpha=0$. Therefore,
$\|f\|_{u,0,1}=\|f^{(1,\dots,1)}_{u}\|_{L^{2}}=\frac{c^{|u|}_{0}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}$
@@
Next, when $\alpha=1$, it follows from the definition (16) that
$\|f\|_{u,1,1}=\sup_{v\subseteq u}\sup_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant 1}^{|u|}\\ \alpha_{j}=1\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}V^{v}_{1}(f^{(\boldsymbol{\alpha}_{u})}_{u})=\sup_{v\subseteq u}V^{v}_{1}(f^{(1,\dots,1)}_{u}).$
Let $g=f^{(1,\dots,1)}_{u}.$ For each $v\subseteq u$, because $((x+1)\exp(x))^{\prime}=(x+2)\exp(x)$,
$V^{(|v|)}_{1}\Bigl{(}g(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})\Bigr{)}=\frac{c^{|u|}_{1}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+2)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|v|/2}\prod_{j\in u\setminus v}(x_{j}+1)\exp(x_{j}).$
Therefore, by equation (12),
$\Big{(}V^{v}_{1}(g)\Big{)}^{2}$
$\leqslant$ $\frac{c^{2|u|}_{1}}{(\prod_{j\in u}j)^{2\gamma}}\Big{(}\int_{0}^{1}(x+2)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|v|}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|-|v|}$
$\leqslant$ $\|f^{(2,\dots,2)}_{u}\|_{L^{2}}^{2},$
while $V^{u}_{1}(g)=\|g^{(1,\dots,1)}\|_{L^{2}}=\|f^{(2,\dots,2)}_{u}\|_{L^{2}}$. Therefore,
$\|f\|_{u,1,1}=\|f^{(2,\dots,2)}_{u}\|_{L^{2}}=\frac{c^{|u|}_{1}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+2)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}$
and we can make $\|f_{s,\gamma,1}\|_{1,1}=1$ by setting
@@
With the above choice of $c_{\alpha}$ for $\alpha=0,1$, we also have $\gamma_{u}=\prod_{j\in}\gamma_{j}$ for $\gamma_{j}=j^{-\gamma}$.
For each integrand $f_{s,\gamma,\alpha}$, we evaluate the performance of the median estimator $\hat{\mu}^{(r)}_{E}$ for $m=1,2,\dots,16$ under two randomization schemes: the completely random design (CRD) and the random linear scrambling (RLS). The estimator $\hat{\mu}^{(r)}_{E}$ is computed as the sample median of $12r-$ 1 independent realizations of $\hat{\mu}_{E}$, where $\hat{\mu}_{E}$ is defined as in equation (2) with precision $E=64$. The observed integration errors in our tested range of $m$ is much larger than $2^{-E}$, ensuring that $\hat{\mu}^{(r)}_{E}$ closely approximates $\hat{\mu}^{(r)}_{\infty}$. Consequently, the error bounds in Theorem 1, 2 and 3 apply to $\hat{\mu}^{(r)}_{E}$ with probability at least $1-2^{-r}$. We set $r=10$, limiting the failure probability per data point to below $0.1\%$. As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the $12r-$ 1 $\hat{\mu}_{E}$ realizations used by RLS. The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [12].
@@
Figure 1: Integration errors for $f_{s,\gamma,\alpha}$ with $\alpha=0$.
@@
Figure 2: Integration errors for $f_{s,\gamma,\alpha}$ with $\alpha=1$.
@@
$\gamma^{\prime}_{u}=\frac{\|f_{s,\gamma,0}\|_{u,1,1}}{\|f_{s,\gamma,0}\|_{1,1}}\leqslant\prod_{j\in u}\Big{(}\frac{c_{0}}{c_{1}}j^{-\gamma}\Big{)}=\prod_{j\in u}\gamma^{\prime}_{j}$
for $\gamma^{\prime}_{j}=(c_{0}/c_{1})j^{-\gamma}$. In general, this analysis extends to arbitrary norms $\|\cdot\|_{\alpha^{\prime},\lambda^{\prime}}$ with $\alpha^{\prime}\in\mathbb{N}_{0},\lambda^{\prime}\in(0,1]$, and the corresponding $\gamma_{j}$ is always proportional to $j^{-\gamma}$. Thus, the decay rate $\gamma$ dominates convergence behavior as $s\to\infty$, irrespective of $\alpha$.
Next, we observe that for a fixed value of $\gamma$, the integration errors exhibit similar behavior as $s$ varies, aligning with our analysis of dimension-independent convergence. Applying equation (44) with parameters $d=1$, $p=\gamma$ and $q=1+\eta$ for an arbitrarily small $\eta>0$, our theory predicts that for large $s$, $\hat{\mu}^{(r)}_{\infty}$ under RLS achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates, provided $\alpha+\lambda+1/2<(\gamma+1)/2$. Empirically, the observed rates approximate $O(2^{-(3/2)m})$ when $\gamma=2$ and $O(2^{-2m})$ when $\gamma=3$. For $\gamma=4$, the convergence rates slightly exceed $O(2^{-2m})$ but remain significantly slower than $O(2^{-(5/2)m})$. This discrepancy is likely due to the limited range of $m$ tested. Notably, while our theory suggests that $\hat{\mu}^{(r)}_{\infty}$ under CRD can attain near $O(2^{-(\alpha+\lambda+1/2)m})$ rates for $\alpha+\lambda+1/2<\gamma$, practical results show that CRD consistently underperforms RLS. We hypothesize that the superior theoretical rates for CRD may only manifest at substantially larger $m$.
Finally, we compare the integration errors under RLS and STD. Across the vast majority of test cases, RLS demonstrates superior performance over STD, reinforcing that the median RQMC approach (via RLS) is a more robust alternative to traditional RQMC (via STD). We also note that STD exhibits convergence rates comparable to those of RLS and substantially surpasses the theoretical $O(2^{-(3/2)m})$ rate derived from the root mean squared error when $\gamma=3$ or $4$. This discrepancy arises because each estimator $\hat{\mu}_{\infty}$ in STD is a special case of $\hat{\mu}^{(r)}_{\infty}$ with $r=1$, which satisfies the rapidly converging probabilistic error bounds in Theorems 2 and 3. However, averaging multiple replicates in STD compromises these bounds, as even a single outlier among the replicates can dominate the error and inflate its magnitude. This is the key mechanism that motivates our shift toward median RQMC approaches.